# Stable-V2A: Synthesis of Synchronized Sound Effects with Temporal and Semantic Controls

### v2

@@ V-C Evaluation Metrics
To perform an objective evaluation of our model, we employ some of the most widely used metrics to attest semantic quality and time alignment in V2A tasks:
TABLE I: Results for Stable-V2A and comparison with other SOTA models on Greatest Hits. Table shows whether the model generates the output conditioned on an audio or text prompt; HRC stands for Human Readable Control and refers to the use of time-varying interpretable signals that sound designers can use to control the generation process (i.e., envelope or onsets). Our model provides the best results, even in the setting of text conditioned generation.
@@ V-C5 Fréchet Audio-Visual Distance
We use E-L1, acc@1, acc@5, acc@10 for the evaluation of the RMS-Mapper, while we use FAD-P, FAD-C, FAD-CL, CLAP-score, FAVD and E-L1 for the evaluation of both Stable-Foley and the complete model, Stable-V2A.
@@ VI-A Complete Model
We compare our complete model, Stable-V2A, with the main V2A models publicly available at the time of writing this paper. More precisely, our baseline models are the aforementioned SpecVQGAN [49], CondFoleyGen [50], Diff-Foley [51], Video-Foley [54] and SyncFusion [30].
@@
As shown in Table I, Stable-V2A achieves improved results across all metrics, including both time alignment and the semantic quality of the generated audio, compared to all baseline models considered. These results highlight the importance of using a strong temporal control, such as the envelope, to guide the generation of audio. Indeed, the second- and third-best models in our tests, Video-Foley and SyncFusion, use envelope and onset tracks for temporal conditioning, providing a stronger guide for temporal alignment than the methods employed by SpecVQGAN, CondFoleyGen, and Diff-Foley.
@@
TABLE II: Results for RMS-Mapper with different RGB and opt. flow encoders.
@@
TABLE III: Results for Stable-Foley for audio prompt and ground-truth RMS with and without frames conditioning.
@@
TABLE IV: Results for Stable-V2A on Walking The Maps.
@@
Stable-V2A 167 0.2556 0.0460 0.6956 3.0682
The use of ControlNet enables us to leverage the prior knowledge of Stable Audio for generating ambient sounds, allowing the production of 44.1 kHz stereo audio effects, matching professional audio production standards. Additionally, the integration of ControlNet with a high $\mathrm{depth\_factor}$ of 20% ensures that our model is both lightweight and fast. In fact, the results we achieve are based on training Stable-Foley on a small dataset of approximately 6 hours, such as Greatest Hits, with a limited number of training steps, thus avoiding the need for substantial time and computational resources.
@@ VI-B1 RMS-Mapper
We also experiment with the feature extractor used in RegNet, that is BN-Inception [83]. This network is trained on images, so the features extracted for one frame are not correlated with those of other frames. However, it is trained on an extremely large dataset (ImageNet [84]), allowing it to generalize well and produce strong results even when encoding video frames, despite not being designed for this purpose. Our experiments show that TC-CLIP achieves better results than BN-Inception, although by a limited margin. This outcome is significant, as TC-CLIP is trained on a smaller dataset (Kinetics-400 [85]) specifically designed for motion recognition. In our opinion, this demonstrates the direction V2A research should pursue: employing models specifically designed to capture temporal context.
Results for RMS-Mapper and relative ablation studies are reported in Table II.
VI-B2 Stable-Foley
we evaluate Stable-Foley by generating audio through ground truth RMS envelopes extracted directly from waveforms. As ablation study, we train the model without the direct conditioning of frames provided by CAVP. The results reported in Table III show that such additional information leads to a perceptible improvement in the semantics of the produced audio, as can be noticed from the metrics for semantic alignment. While there is no perceivable enhancement in temporal alignment; this result shows that temporal conditioning is contributed entirely by the ControlNet input, demonstrating the strength of such conditioning, which is the desired result.
@@ VII Conclusion and Discussion
In this paper, we present a model for generating an audio track semantically and temporally aligned to a silent input video, called Stable-V2A. The model is divided into two distinct parts, trained separately and joined only at inference time: RMS-Mapper, which maps a representative envelope of the audio to be generated directly from the reference video, and Stable-Foley, a sound synthesis model that, through the use of the predicted envelope and other semantic controls, generates the final output of the model. Our sound synthesis model is based on Stable Audio Open and, to the best of our knowledge, this is the first time that such a state-of-the-art model for audio generation is used in the context of Video-to-Audio.
Although the audio synthesis model tracks with high accuracy the envelope used as a control, mapping that envelope from the video represents the bottleneck of our model, still leaving room for many potential research works.
@@ VIII Acknowledgements
E.P. and L.C. were partially supported by PRIN 2022 project no. 2022AL45R2 (EYE-FI.AI, CUP H53D2300350-0001.

### v3

@@ V-C Evaluation Metrics
To perform an objective evaluation of our model, we employ some of the most widely used metrics to attest semantic quality and time alignment in V2A tasks. We use E-L1, acc@1, acc@5, acc@10 for the evaluation of the the video model, while we use FAD-P, FAD-C, FAD-CL, CLAP-score, FAVD and E-L1 for the evaluation of both the audio synthesis model and the complete model, FOL·AI.
TABLE I: Results for FOL·AI and comparison with other SOTA models on Greatest Hits. Table shows whether the model generates the output conditioned on an audio or text prompt; HRC stands for Human Readable Control and refers to the use of time-varying interpretable signals that sound designers can use to control the generation process (i.e., envelope or onsets). Our model provides the best results, even in the setting of text conditioned generation.
@@ VI-A Complete Model
We compare our complete model, FOL·AI, with the main V2A models publicly available at the time of writing this paper. More precisely, our baseline models are the aforementioned SpecVQGAN [35], CondFoleyGen [36], Diff-Foley [37], Video-Foley [42] and SyncFusion [38].
@@
As shown in Table I, FOL·AI achieves improved results across all metrics, including both time alignment and the semantic quality of the generated audio, compared to all baseline models considered. These results highlight the importance of using a strong temporal control, such as the envelope, to guide the generation of audio. Indeed, the second- and third-best models in our tests, Video-Foley and SyncFusion, use envelope and onset tracks for temporal conditioning, providing a stronger guide for temporal alignment than the methods employed by SpecVQGAN, CondFoleyGen, and Diff-Foley.
@@
TABLE II: Results for the video model with different RGB and opt. flow encoders.
@@
TABLE III: Results for the audio synthesis model for audio prompt and ground-truth RMS with and without frames conditioning.
@@
TABLE IV: Results for FOL·AI on Walking The Maps.
@@
FOL·AI 167 0.2556 0.0460 0.6956 3.0682
The use of ControlNet enables us to leverage the prior knowledge of Stable Audio for generating ambient sounds, allowing the production of 44.1 kHz stereo audio effects, matching professional audio production standards. In fact, compared with the baseline methods reported, ours is the only one that generates audio at a standard rate for professional audio. Additionally, the integration of ControlNet with a high $\mathrm{depth\_factor}$ of 20% ensures that our model is both lightweight and fast. Indeed, the results we achieve are based on training the model on a small dataset of approximately 6 hours, such as Greatest Hits, with a limited number of training steps, thus avoiding the need for substantial time and computational resources.
@@ VI-B1 Different encoders in the video model
We also experiment with the feature extractor used in RegNet, that is BN-Inception [73]. This network is trained on images, so the features extracted for one frame are not correlated with those of other frames. However, it is trained on an extremely large dataset (ImageNet [74]), allowing it to generalize well and produce strong results even when encoding video frames, despite not being designed for this purpose. Our experiments show that TC-CLIP achieves better results than BN-Inception. This outcome is significant, as TC-CLIP is trained on a smaller dataset (Kinetics-400 [75]) specifically designed for motion recognition. In our opinion, this demonstrates the direction V2A research should pursue: employing models specifically designed to capture temporal context.
Results for the video model and related ablation studies are reported in Table II.
VI-B2 Use of video features in the audio synthesis model
We evaluate the audio synthesis model by generating audio through ground truth RMS envelopes extracted directly from waveforms. As ablation study, we train the model without the direct conditioning of frames provided by CAVP. The results reported in Table III show that such additional information leads to a perceptible improvement in the semantics of the produced audio, as can be noticed from the metrics for semantic alignment. While there is no perceivable enhancement in temporal alignment; this result shows that temporal conditioning is contributed entirely by the ControlNet input, demonstrating the strength of such conditioning, which is the desired result.
@@ VII Conclusion and Discussion
In this paper, we present a novel model for generating an audio track semantically and temporally aligned to a silent input video, called FOL·AI. The model is divided into two distinct parts, trained separately and joined only at inference time: the video model, which maps a representative envelope of the audio to be generated directly from the reference video, and the audio synthesis model that, through the use of the predicted envelope and other semantic controls, generates the final output of the model. Our sound synthesis model leverages Stable Audio Open and, to the best of our knowledge, this is the first time that such a state-of-the-art model for audio generation is used in the context of Video-to-Audio. The main contributions of FOL·AI also include the novel application of ControlNet to Stable Audio’s DiT architecture for precise temporal control using RMS envelopes, the introduction of the Walking The Maps dataset focused on the challenging Foley task of footstep synthesis, and achieving state-of-the-art performance in V2A tasks while generating high-fidelity 44.1 kHz stereo audio.
Although the audio synthesis model tracks with high accuracy the envelope used as a control, mapping that envelope from the video represents the bottleneck of our model, still leaving room for potential future research works.
@@ VIII Acknowledgements
E.P. and L.C. were partially supported by PRIN 2022 project no. 2022AL45R2 (EYE-FI.AI, CUP H53D2300350-0001).